OpenStudy (anonymous):

solve the equation by finding square roots 3x^2=21 but how do you do the work?

OpenStudy (anonymous):

We need to get x by itself, what do you get when you divide both sides by 3?

OpenStudy (anonymous):

7...then square root the 7 so its +or- the square root of 7??

OpenStudy (anonymous):

yes, nice job!

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

wanna help again? factor 4x^2+28x-32=0

OpenStudy (anonymous):

if i may: we can begin by factoring out a 4, giving us: 4(x^2+7x-8)=0 Now we have but to find two number that multiply to give us -8 and add to give us 7, we discern that those two numbers are -1 and 8, so the quadratic then factors to: 4(x-1)(x+8)=0 Now we set each term to 0: 4=0 x-1=0--->x=1 x+8=0--->x=-8

OpenStudy (anonymous):

i get it thanks so much